Within 11 years, the age of peter will be half the square of the age he was 13 years ago. Calculate current age of peter.
Answer by stanbon(75887)

Within 11 years, the age of peter will be half the square of the age he was 13 years ago. Calculate current age of peter.
-----------
Equation:
p+11 = (1/2)(p-13)^2
--------
2p + 22 = p^2 - 26p + 169
--------
p^2 - 28p + 147 = 0
-------
peter's age = 21
---------------
Cheers,
Stan H.
-------------